17. From the following joint probability distribution of X and Y,f (a) P(X= Y], P [X > Y] (b) the distribution of X+ Y (c) E(X), E(Y), Var(X), Var(Y) (d) Cov(X, Y), Corr(X, Y) nd: 0 .3 .25 .05 .2 19 T
Answer #1
X \ Y 0 1 2 P[X=x]
0 0.1 0.3 0.05 0.45
1 0.2 0.25 0.1 0.55
P[Y=y] 0.3 0.55 0.15 1

a) P[X= Y] = P[X-Y = 0]

= P[X=0,Y=0]+P[X=1,Y=1 ]

= 0.1 +0.25

= 0.35

b) P[X>Y ] = P[X=1 , Y = 0] = 0.2

c) E[X]= \Sigma \left xP[X=x] =0 \times 0.45 + 1 \times 0.55

So, E[X] = 0.55

  Elx2l Σ x2PlX x-02 x 0.45 + 12 x 0.55 = 0.55

So, Var(X) = E[X2] - E2[X] = 0.55 - 0.552 = 0.2475

  

  E[Y]= \Sigma \left yP[Y=y] =0 \times 0.3 + 1 \times 0.55+ 2 \times 0.15

So, E[Y] = 0.85

E[Y^2]= \Sigma \left y^2P[Y=y] =0^2 \times 0.3 + 1^2 \times 0.55 + 2^2 \times 0.15= 1.15

So, Var (Y) = 1.15 - 0.852 = 0.4275  

d) Again   Cov(X,Y) = E(XY)-E(X)E(Y)

now,   E(XY) = \Sigma \left xy P[X=x, Y=y]

= 0+ (1\times 1\times 0.25) + (1\times2\times 0.1) = 0.45

So, Cov(XY) = 0.45 - 0.55\times0.85 = - 0.0175

and Corr(X,Y) = \frac{Cov(X,Y)}{\sqrt{Var(X)Var(Y)}}

Corr(X,Y) = \frac{-0.0175}{\sqrt{0.2475 \times 0.4275}} = -0.0538

If you have any doubts, ask me below in the comment section. Thank you